# Assessing and improving reliability of neighbor embedding methods: a map-continuity perspective

### v1

@@ 2.7 Use case 2: enhancing interpretation of single-cell data
The distribution of large singular values provides an informative indicator of the presence of spurious sub-clusters, which echoes the irregular landscape of the LOO loss function (Supplementary Figure S3a, c). We verified the distribution difference extensively using multiple statistical tests including Spearman’s rank correlation tests between singularity scores and distances to cluster centers, F-tests and permutation tests for a local regression model (singularity scores regressed against locations). We find that the $p$-values for all tests drop to a significant level after appropriately choosing the perplexity (Methods A.3.5). First, in the Spearman’s Rank correlation tests, at low perplexities, the correlations between singularity scores and locations are not significant in all five clusters in the mESCs data and all six clusters in the mouse brain ATAC data, with respective averaged $p$-values being $0.36$ and $0.88$ (after Holm-Bonferroni correction for multiple testing) (Supplementary Table S4). In contrast, after increasing the perplexities to the elbow points of the singularity score plots (Figure 6b,d), we observed significant correlations (at a significance level of $0.05$) in four of five clusters in the mESCs data and three of six clusters in the mouse brain ATAC data. Next, for the F-tests and permutation tests for the local regression model, we tested whether the slope parameters in the local regression model are zero, and observed that $p$-values of both tests dropped from around $0.2$ to less than $10^{-14}$ in the mESCs dataset and from around $0.09$ to less than $10^{-15}$ in the mouse brain ATAC dataset, which indicated significant dependence of singularity scores on locations after increasing perplexity to the elbow points. The tests show that embedding points with high singularity scores transit from a disordered distribution at a low perplexity to an ordered distribution along the edges of clusters at a higher perplexity. Indeed, this location dependence can be explained by the geometry of LOO loss landscapes—a small perplexity may lead to many local minima in visually random locations (Supplementary Figure S3a, c) which then result in spurious sub-clusters, while a larger perplexity leads to more regular loss landscapes (Supplementary Figure S3b, d).
We also observe that the degree of FI discontinuity, as indicated by the magnitude of the singularity scores, decreases rapidly until the perplexity reaches the elbow point (Figure 6b, d). If we choose a perplexity larger than the elbow point, the spurious sub-clusters largely disappear from embeddings. This disappearance of sub-clusters is aligned with the improvement of neighborhood preservation (Figure 6b, d), as measured by the correlation of nearest-neighbor distances between the input space and the embedding space (Methods A.3.3). However, we would not suggest increasing perplexity excessively, as it may cause clusters to merge [25] and result in the loss of certain genuine local or microscopic structures [6]. Also, a higher perplexity often leads to longer computational running time [26]. Therefore, we suggest choosing perplexity around the elbow point.
2.8 Computational costs
@@
The computational complexity for calculating the perturbation score per point is $O(n^{2})$ flops, where the theoretical bottleneck lies in the calculation of similarity scores $(v_{i,j})_{i\leq j}$. Practically, our LOO loss has significantly sped up the evaluation of embedding points under perturbation by reducing the original optimization loss with $O(n^{2})$ terms and $O(n)$ variables to $O(n)$ terms and $O(1)$ variables. Overall, it takes 35.2 seconds on a MacBook Air with Apple M2 chip to calculate the exact perturbation score per point for CIFAR-10 images in Figure 1.
We provide two optional methods in our R package as practical approximations to improve computation speed. First, we offer an approximation method which reduces the time of calculating perturbation scores to $7.1$ seconds per point by reusing certain pre-computed quantities for approximation (Methods A.2.1). Second, we implement a pre-screening step to select a subset of the embedding points as the most likely high-scoring points, instead of calculating the scores for every point. We use the heuristics that high perturbation scores are often found at the peripheries of clusters. Such pre-screening of data points can significantly reduce computational effort while still providing a comparable assessment of OI discontinuity locations across the entire embedding (Supplementary Figure S5). Combining the two options leads to a reduction of running time by roughly 97%. In addition, since our perturbation scores are defined on a point-wise basis, they are highly parallelizable and can be further accelerated in computation.
@@
The computational complexity for calculating the singularity scores for the entire dataset is $O(n^{2})$ flops, primarily due to the matrix operations when calculating Hessian matrices.
Notably, the computation of singularity scores is relatively fast, requiring only 15.9 seconds in total on a MacBook Air with Apple M2 chip for a dataset with 5,000 points (Figure 1d). This fast speed is a desirable feature for tuning hyperparameters across multiple candidates. In addition, we expect further speedup in future improvements by exploiting the sparsity of similarity matrix at a low perplexity.
2.9 Comparison with other assessment metrics
@@
There are multiple recent papers on assessing and improving the reliability of neighbor embedding methods. None of these papers view the observed artifacts as an intrinsic map discontinuity, and as a result, cannot reliably identify topological changes in their proposed diagnosis. For illustration, here we compare our method with EMBEDR [28], scDEED [27], and DynamicViz [29].
@@
  • EMBEDR identifies dubious points in an embedding by using statistical significance estimates as point-wise reliability scores. This process begins by computing point-wise KL divergences between the kernels in the input space and those in the embedding space, followed by a permutation test to determine whether the neighborhood preservation is significantly better than random chance. Lower $p$-values from the test indicate higher embedding reliability.
@@
  • scDEED calculates point-wise $p$-values by conducting a similar permutation test on the correlations of nearest-neighbor distances, in replace of KL divergences. Similarly, lower $p$-values indicate higher embedding reliability.
@@
  • DynamicViz employs a bootstrap approach to assess the stability of embeddings. A point-wise variance score is constructed based on resampling, defined as the mean variance of distances to the neighbors. Embedding points with lower variance scores are considered more reliable.
Both EMBEDR and scDEED focus on the faithfulness of neighborhood preservation. They treat the neighborhood structure of each point as a signal and test it against a randomized noisy baseline. However, neither methods aim to detect issues of topological changes and thus cannot accurately locate the discontinuity points. In contrast, our perturbation scores are designed to capture embedding points that are close to the intrinsic discontinuity of the embedding maps. In this regard, perturbation scores are better at locating the changes of global structures.
We illustrate the advantages of our perturbation scores in identifying topological changes on a simulated Swiss roll dataset, where the t-SNE algorithm erroneously splits the smooth Swiss roll manifold into two disconnected pieces (Figure 7a). This visualization artifact represents a severe qualitative change: a smooth connected manifold is visualized as disconnected components, as reflected by the OI discontinuity. The perturbation scores unambiguously highlight the unreliable points exactly at the location where the disconnection occurs (Figure 7b). In comparison, both EMBEDR and scDEED evaluate most of the embedding points as unreliable and dubious, failing to identify the exact location of the discontinuity (Figure 7c, d).
In addition, while DynamicViz can roughly identify the location of the discontinuity (Figure 7e), it still fails to pinpoint the exact points where the split occurs due to the inherent variance in the bootstrap approach. Using a different dataset, we identify another limitation of DynamicViz. Specifically, the bootstrap procedure fails to robustly characterize the discontinuity locations in areas with lower point density, as these areas are prone to insufficient sampling (Supplementary Figure S7). In contrast, our perturbation scores are less susceptible to the low density regions.
@@
Figure 7: Comparing perturbation scores with three diagnostic scores for the t-SNE embedding on the simulated Swiss roll dataset. a t-SNE embedding of $n=1000$ simulated points from the Swiss roll manifold. The colors correspond to the ground-truth spiral angles of the points. t-SNE algorithm erroneously breaks the smooth manifold into two disconnected parts, which indicates OI discontinuity. b Perturbation scores clearly mark the unreliable embedding points where disconnection (discontinuity) occurs. c EMBEDR suggests that most embedding points are unreliable (lower $p$-values are more reliable), but it does not identify the discontinuity location. d ScDEED evaluates most embedding points as dubious, but similar to EMBEDR, it does not identify the discontinuity location. e DynamicViz marks both the discontinuity location and the areas at both ends of the Swiss roll as unstable, making it difficult to distinguish the actual discontinuity locations. Furthermore, while it can roughly identify the discontinuity location, it still fails to pinpoint the exact points where the split occurs.
2.10 Theoretical insights: landscape of LOO loss
@@
By analyzing the LOO loss function in Eqn. 3 under a simple setting, we will show that OI discontinuity is caused by a hyperbolic saddle point in the LOO loss function, thereby theoretically justifying Figure 3b.
Suppose that $n$ input points ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$ are generated from a data mixture with two well-separated and balanced groups, where the first group is represented by the index set ${\mathcal{I}}_{+}\subset\{1,2,\ldots,n\}$ with $|{\mathcal{I}}_{+}|=n/2$ and the second group represented by ${\mathcal{I}}_{-}=\{1,2,\ldots,n\}\setminus{\mathcal{I}}_{+}$. Without loss of generality, we assume that the mean vectors of $({\bm{y}}_{i})_{i\in{\mathcal{I}}_{+}}$ and $({\bm{y}}_{i})_{i\in{\mathcal{I}}_{-}}$ are $\bm{\theta}$} and $\bm{\theta}$} respectively, since embeddings are invariant to global shifts and rotations. Equivalently, we write
${\bm{y}}_{i}=\begin{cases}{\bm{\theta}}+\bm{\delta}_{i}&i\in{\mathcal{I}}_{+}\\ -{\bm{\theta}}+\bm{\delta}_{i}&i\in{\mathcal{I}}_{-}\end{cases}$
@@
where $\sum_{i\in{\mathcal{I}}_{+}}\bm{\delta}_{i}=\sum_{i\in{\mathcal{I}}_{-}}\bm{\delta}_{i}=\mathbf{0}$. To simplify the loss function, we make an asymptotic assumption: consider (implicitly) a sequence of problems where input data have increasing distances between the two groups, so we expect increasing separation of clusters in the embedding space:
$\|{\bm{\theta}}\|\to\infty,\qquad\max_{i\leq n}\|\bm{\delta}_{i}\|=O(1)\,.$
@@
Now considering adding an input point (‘mixed’ point) to a location close to the midpoint of the two groups. We assume that its similarity to the other inputs is
@@
for $1\leq i\leq n$, where $p_{0}>0$ and $\varepsilon$ is a small perturbation parameter. This assumption is reasonable because the similarity of the added point ${\bm{x}}={\bm{x}}^{\varepsilon}$ has roughly equal similarities to existing inputs up to a small perturbation. We make the asymptotic assumption $\|{\bm{\theta}}\|^{-1}\asymp\varepsilon$, namely $\varepsilon\|{\bm{\theta}}\|=O(1)$ and $\varepsilon\|{\bm{\theta}}\|]^{-1}=O(1)$.
@@
Consider the LOO loss function for t-SNE given in Eqn. 3 and 2. Under the above stated assumptions, the negative gradient of the loss is
@@
where ${\bm{y}}_{\mathbin{\!/\mkern-5.0mu/\!}}={\bm{\theta}}{\bm{\theta}}^{\top}{\bm{y}}/\|{\bm{\theta}}\|^{2}$ is projection of $\bm{y}$} in the direction of $\bm{\theta}$}, and ${\bm{y}}_{\bot}={\bm{y}}-{\bm{y}}_{\mathbin{\!/\mkern-5.0mu/\!}}$.
@@
  • The hyperbolic term indicates the unstable saddle point of the loss at ${\bm{y}}=\mathbf{0}$. Indeed, it is exactly the tangent vector of a hyperbola, so in the embedding force (negative gradient) field there is a pull force towards the x-axis and a push force away from the y-axis (Figure 8).
  • The perturbation term reflects the effects of input point ${\bm{x}}^{\varepsilon}$. It tilts the negative gradients slightly in the direction of $\bm{\theta}$} if $\varepsilon>0$ or $\bm{\theta}$} if $\varepsilon<0$, which causes the algorithm to jump between widely separated local minima of $L({\bm{y}};{\bm{x}})$ under small perturbations.
Figure 8: Negative gradient fields of the real/theoretical LOO loss. a We draw the negative gradient fields (force fields) $\nabla_{{\bm{y}}}L({\bm{y}};{\bm{x}}^{\varepsilon})$ based on the LOO loss under the same setting as in Figure 3b. b We draw a similar field plot based on the hyperbolic term $\frac{{\bm{y}}_{\mathbin{\!/\mkern-5.0mu/\!}}-\bm{y}_{\bot}}{\|{\bm{\theta}}\|^{2}}$ from Theorem 1, where we take ${\bm{\theta}}=({\bm{c}}_{1}-{\bm{c}}_{2})/2$ and ${\bm{c}}_{1},{\bm{c}}_{2}$ are the centers of two clusters in the embedding. In addition, we add loss contours to both plots, which show hyperbolic paraboloids around the origin. We observe excellent alignment between the negative gradient field of the LOO loss and that of the theoretical analysis. Both field plots show a pull force towards the x-axis and a push force away from the y-axis.
3 Discussion
In this paper, we developed a new framework for interpreting distortions in neighbor embedding methods as instances of map discontinuity by leveraging the LOO idea. Through our proposed LOO-map, we introduced two diagnostic scores aimed at identifying and addressing two distinct types of map discontinuities—OI and FI discontinuities. While our method has shown promising results, we acknowledge that our work may not account for all types of distortions. For example, factors such as initialization schemes, the choices of iterative algorithms, and other hyperparameters, may introduce different types of distortions. Additionally, we recognize the absence of a formal mathematical framework for rigorously characterizing the LOO-map.
In future research, we aim to establish connections between classical parametric embedding maps and the implicit embedding maps from neighbor embedding methods, which may fully address the topological issues and enhance both interpretability and faithfulness. One possible theoretical approach is [14]. Moreover, while our approach performs well on moderately sized datasets, further work is needed to develop scalable version of these methods. This includes developing more efficient optimization implementation, leveraging sparsity and tree-based approximation, and fully exploiting parallelism in computation.
Appendix A Methods
A.1 Verify Leave-one-out assumption empirically
@@
Let ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{n}]^{\top}$ be the input data matrix, and ${\bm{Y}}=[{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}$ be the matrix of embedding points. We then add one point $\bm{x}$} to $\bm{X}$} to have the new input data ${\bm{X}}_{+}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{n},{\bm{x}}]^{\top}$. We then run t-SNE algorithm to obtain the embedding of $\bm{X}$$_{+}$ as $\widetilde{\bm{y}}_{1},\ldots,\widetilde{\bm{y}}_{n},\widetilde{\bm{y}}$$^{\top}$. Denoted $\widetilde{\bm{y}}_{1},\ldots,\widetilde{\bm{y}}_{n}]$ as $\widetilde{{\bm{Y}}}$. To verify LOO empirically, we keep track of the difference between $\bm{Y}$} and $\widetilde{{\bm{Y}}}$:
$\epsilon_{n}=\frac{1}{\|{\bm{Y}}\|_{F}}\|{\bm{Y}}-\widetilde{\bm{Y}}\|_{F}$
@@
We initialize the t-SNE algorithm in the second run by the embedding points we obtain from the first run: when calculating the embedding of $\bm{X}$$_{+}$, we use $\bm{Y}$} as the initialization for the first $n$ points. This initialization scheme aims to address two issues: (i) the loss function in a neighbor embedding method is invariant to a global rotation and a global shift of all embedding points, so it is reasonable to choose embedding points with an appropriate initialization. (ii) There are potentially multiple local minima of the loss function due to non-convexity. We verify the LOO assumption at a given local minimum (namely $\bm{Y}$}) obtained from the first run.
@@
The experiment is conducted with different sample size $n$ and with different types of dataset (simulated cluster data, simulated manifold data, real single-cell data, deep learning feature data). The comprehensive results showing the values of $\epsilon_{n}$ under different settings are presented in Supplementary Table S2. We observe that the approximation errors $\epsilon_{n}$ are small and generally decreasing in $n$, which supports our LOO assumption.
A.2 Two diagnosis scores
@@
For implementation convenience, our calculation of the two diagnosis scores are based on modifying an input point instead of adding a new input point. According to the LOO assumption, the difference is negligible.
A.2.1 Perturbation score
@@
Given an input data matrix ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{n}]^{\top}$ and its embedding matrix ${\bm{Y}}=[{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}$, we view ${\bm{y}}_{i}$ as the mapping of ${\bm{x}}_{i}$ by the partial LOO-map ${\bm{f}}_{i}$:
${\bm{f}}_{i}({\bm{x}})={\rm argmin}_{{\bm{y}}\in\mathbb{R}^{2}}L_{i}({\bm{y}};{\bm{x}}),\qquad\text{where}$ (5)
$L_{i}({\bm{y}};{\bm{x}})=\sum_{k\neq i}\mathcal{L}\Big(w({\bm{y}}_{k},{\bm{y}});v_{i,k}(\bar{\bm{X}})\Big)+Z(\bar{\bm{Y}}).$
@@
where $\bar{\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{i-1},{\bm{x}},{\bm{x}}_{i+1},\ldots,{\bm{x}}_{n}]^{\top}$ differs from $\bm{X}$} only at the $i$-th input point, and $\bar{\bm{Y}}=[{\bm{y}}_{1},\ldots,{\bm{y}}_{i-1},{\bm{y}},{\bm{y}}_{i+1},\ldots,{\bm{y}}_{n}]^{\top}$ has frozen embedding points except for the $i$-th point which is the decision variable in the optimization problem. This partial LOO-map ${\bm{f}}_{i}$ is based on perturbing (or modifying) a single input point rather than adding a new point, thus maintaining $n$ points in total. According to the LOO assumption, ${\bm{f}}_{i}\approx{\bm{f}}$, so we calculate perturbation score for the $i$-th point based on ${\bm{f}}_{i}$.
@@
To assess the susceptibility of ${\bm{y}}_{i}$ under moderate perturbation in ${\bm{x}}_{i}$, we apply a perturbation of length $\lambda$ in the direction of $\bm{e}$} to ${\bm{x}}_{i}$ and measure the resulting change in mapping determined by the partial LOO-map ${\bm{f}}_{i}$. When implementing, we search the perturbation directions among the first 3 principal directions of the data $\{{\bm{e}}_{1},{\bm{e}}_{2},{\bm{e}}_{3}\}$ and their opposites $\{-{\bm{e}}_{1},-{\bm{e}}_{2},-{\bm{e}}_{3}\}$, and the perturbation length $\lambda$ is specified by the user. In this way, we can define the perturbation score of the $i$-th data point as
$\max_{{\bm{e}}\in\{\pm{\bm{e}}_{1},\pm{\bm{e}}_{2},\pm{\bm{e}}_{3}\}}\|{\bm{f}}_{i}({\bm{x}}_{i}+\lambda{\bm{e}})-{\bm{y}}_{i}\|_{2}.$ (6)

### v2

@@ Results
The distribution of large singular values indicates spurious sub-clusters, reflecting the irregular LOO loss landscape (Supplementary Fig. S3a, c). We extensively validated the distribution difference between small and large perplexities through statistical tests, including Spearman’s correlation test between singularity scores and cluster center distances, F-tests, and permutation tests for local regression models (singularity scores regressed against locations). At low perplexities, Spearman’s correlation tests showed non-significant results for all five mESCs clusters and five mammary epithelial cell clusters (average $p$-values: 0.36 and 0.29, Supplementary Table S5). Increasing perplexities to the singularity score elbow points (Fig. 6b, d) yielded significant correlations in four of five mESCs clusters and five of eight mammary epithelial cell classes. Similarly, F-tests and permutation tests showed $p$-values dropping from $\sim 0.3$ to <0.001<0.001 (mESCs) and $10^{-13}$ (mammary epithelial cells), confirming the dependence of singularity scores on location at higher perplexities. This transition aligns with LOO loss geometry: low perplexities create scattered local minima, forming spurious sub-clusters (Supplementary Fig. S3a, c), whereas higher perplexities smooth the loss landscape (Supplementary Fig. S3b, d), reducing artifacts.
We also observe that the degree of FI discontinuity, as indicated by the magnitude of the singularity scores, decreases rapidly until the perplexity reaches the elbow point (Fig. 6b, d). Beyond the elbow point, the spurious sub-clusters largely disappear, aligning with the improvement of neighborhood preservation (Fig. 6b, d), as measured by the nearest-neighbor distance correlation between the input and embedding spaces (Methods). However, we would not suggest increasing perplexity excessively, as it may merge clusters [26], result in the loss of certain microscopic structures [3], and often lead to longer computational running time [27]. Therefore, we suggest choosing a perplexity around the elbow point.
Computational costs
@@
Theoretically, the computational complexity for solving the LOO loss optimization takes $O(n)$ flops, instead of $O(n^{2})$ flops of the original loss which involves every pairwise interaction term. Practically, our R package has the following running time.
  • For exact perturbation scores, it takes 35.2 seconds to compute the score per point for the CIFAR-10 images in Fig. 1 on a MacBook Air (Apple M2 chip).
  • Leveraging pre-computed quantities, we also provide an approximation method to reduce the running time per point to 7.1 seconds, while preserving high accuracy relative to the exact score (Supplementary Fig. S8).
  • In addition to the approximation, we introduce a pre-screening step to increase the computational efficiency by 14X for the same dataset. This pre-screening step identifies a subset of embedding points most likely to yield high scores, and thus significantly reduces computational cost while still providing a comparable assessment of OI discontinuity locations (Supplementary Fig. S7). Combining the approximation method and the pre-screening step results in an average of 0.47 seconds.
@@
Theoretically, the computational complexity for calculating the singularity scores for the entire dataset is $O(n^{2})$ flops, primarily due to matrix operations when calculating Hessian matrices. Practically, the running time for computing singularity scores for CIFAR-10 is 15.9 seconds for all 5,000 points on a MacBook Air (Apple M2 chip).
Comparison with other assessment metrics
@@
There are multiple recent papers on assessing and improving the reliability of neighbor embedding methods. None of these papers view the observed artifacts as an intrinsic map discontinuity, and as a result, cannot reliably identify topological changes in their proposed diagnosis. For illustration, we compare our method with EMBEDR [29], scDEED [28], and DynamicViz [30].
@@
  • EMBEDR identifies dubious embedding points by using statistical significance estimates as point-wise reliability scores. This process begins by computing point-wise KL divergences between the kernels in the input and embedding spaces, followed by a permutation test to determine whether the neighborhood preservation is significantly better than random chance. Lower $p$-values from the test indicate higher embedding reliability. EMBEDR selects the perplexity by minimizing the median $p$-values.
@@
  • scDEED calculates point-wise $p$-values by conducting a similar permutation test on the correlations of nearest-neighbor distances. Similarly, lower $p$-values indicate higher embedding reliability. ScDEED provides two approaches for parameter selection based on dubious embedding points: the first locates the elbow point and the second selects the perplexity to minimize the number of dubious points.
@@
  • DynamicViz employs a bootstrap approach to assess the stability of embeddings. Point-wise variance scores are constructed based on resampling, defined as the average variance of distances to the neighbors. Embedding points with lower variance scores are considered more reliable. DynamicViz selects the perplexity by minimizing the median variance score.
Detecting distortion of global structure.
Compared with existing methods, our perturbation scores have the following advantages. First, perturbation scores are better at locating the topological changes of global structures by pinpointing the exact points. By design, they capture embedding points close to the intrinsic discontinuity of the embedding map. In a simulated Swiss roll dataset, t-SNE erroneously splits the smooth manifold into two disconnected pieces (Fig. 7a), which is a severe visualization artifact caused by OI discontinuity. Our perturbation scores accurately highlight unreliable points exactly at the disconnection location (Fig. 7b). In contrast, EMBEDR and scDEED label most points as unreliable, failing to pinpoint the discontinuity (Fig. 7c, d), as they emphasize neighborhood preservation rather than topological changes. DynamicViz identifies the general region but lacks precision (Fig. 7e).
Second, perturbation scores are also robust to low-density regions. In the simulated Gaussian mixture dataset (Supplementary Fig. S9a), DynamicViz fails to accurately characterize discontinuity locations in areas with a lower point density, as these areas are prone to insufficient sampling (Supplementary Fig. S9c). In contrast, our perturbation scores are more robust to the low-density regions (Supplementary Fig. S9b).
Aiding hyperparameter selection.
Compared with existing methods, our singularity score consistently selects a perplexity that is neither too small nor too large, thus reducing sub-clusters yet still producing fine-grained structures. We illustrate the advantage of consistency using three datasets (Supplementary Table S7).
For the mouse embryonic cell differentiation data (Fig. 6a), scDEED recommends two approaches for perplexity selection; the first is based on the elbow point and yields $3$, and the second is based on minimizing the number of dubious points and does not produce a unique value (Supplementary Fig. S10a). EMBEDR fails to suggest a valid hyperparameter because we encountered errors potentially due to a small dataset size. DynamicViz and singular scores select moderate perplexity (20 and 25), reducing spurious sub-clusters compared to perplexity of 3 (Supplementary Fig. S10b) and achieving the higher neighborhood preservation score (0.5594 (singularity score, highest), 0.4955 (scDEED), 0.5524 (DynamicViz)).
For the mouse brain chromatin accessibility data (Supplementary Fig. S5a), scDEED selects 10 (elbow point) and 145 (minimizing dubious points). EMBEDR chooses perplexity of 145, showing a tendency of favoring larger perplexity that is also observed by [28]. DynamicViz selects perplexity of 10. Our singularity score selects perplexity of 95 (Supplementary Fig. S11a). By visual inspection, perplexity of 10 is inappropriately small because visualization exhibits numerous spurious sub-clusters. In contrast, perplexities of 95 and 145 avoid spurious sub-clusters while maintaining fine-grained structures (Supplementary Fig. S11b). Quantitatively, the perplexities suggested by singular scores, scDEED and EMBEDR lead to similar neighborhood preservation scores (0.4108, 0.4223, 0.4223).
In the mouse mammary epithelial cell dataset, similar phenomena are observed: our singularity score selects a balanced perplexity while scDEED and EMBEDR select perplexities that are either too small or too large, and DynamicViz lacks scalability for large datasets due to its bootstrap-based approach, which requires repeated execution of visualization algorithms (Supplementary Fig. S12). Overall, the singularity score offers robust guardrail perplexities that significantly reduce spurious sub-clusters while producing informative visualization.
@@
Figure 7: Comparing perturbation scores with three diagnostic scores for the t-SNE embedding on the simulated Swiss roll dataset. a The t-SNE embedding of $n=1000$ simulated points from the Swiss roll manifold under perplexity $150$. The colors correspond to the ground-truth spiral angles of the points. t-SNE algorithm erroneously breaks the smooth manifold into two disconnected parts, which indicates OI discontinuity. b Perturbation scores clearly mark the unreliable embedding points where disconnection (discontinuity) occurs. c EMBEDR suggests that most embedding points are unreliable (lower $p$-values are more reliable), but it does not identify the discontinuity location. d ScDEED evaluates most embedding points as dubious, but similar to EMBEDR, it does not identify the discontinuity location. e DynamicViz marks both the discontinuity location and the areas at both ends of the Swiss roll as unstable, making it difficult to distinguish the actual discontinuity locations. Furthermore, while it can roughly identify the discontinuity location, it still fails to pinpoint the exact points where the split occurs.
Theoretical insights: landscape of LOO loss
@@
By analyzing the LOO loss function in Equation 3 under a simple setting, we will show that OI discontinuity is caused by a hyperbolic saddle point in the LOO loss function, thereby theoretically justifying Fig. 3b.
[…]
@@
where $\sum_{i\in{\mathcal{I}}_{+}}{\mathbf{\updelta}}_{i}=\sum_{i\in{\mathcal{I}}_{-}}{\mathbf{\updelta}}_{i}=\mathbf{0}$. To simplify the loss function, we make an asymptotic assumption: consider (implicitly) a sequence of problems where input data have increasing distances between the two groups, so we expect an increasing separation of clusters in the embedding space:
[…]
@@
Now consider adding an input point (‘mixed’ point) to a location close to the midpoint of the two groups. We assume that its similarity to the other inputs is
@@
[…]
@@
Consider the LOO loss function for t-SNE given in Equation 2 and 3. Under the assumptions stated above, the negative gradient of the loss is
@@
[…]
@@
  • The hyperbolic term indicates the unstable saddle point of the loss at ${\mathbf{y}}=\mathbf{0}$. Indeed, it is exactly the tangent vector of a hyperbola, so in the embedding force (negative gradient) field there is a pull force towards the x-axis and a push force away from the y-axis (Fig. 8).
[…]
Discussion
We developed a framework to interpret distortions in neighbor embedding methods as map discontinuities by leveraging the LOO strategy. Based on our LOO-map, we introduce two diagnostic scores to identify OI and FI discontinuities. While being generally effective, our method may not capture all distortion patterns, as factors like initialization, iterative algorithms, and other hyperparameters can introduce different types of distortions. We also recognize the absence of a formal mathematical framework for rigorously characterizing the LOO-map.
In future research, we aim to explore links between classical parametric and implicit embedding maps to fully address topological issues and improve interpretability. We also aim to enhance the scalability of our methods through efficient optimization, sparsity, tree-based approximations, and parallel computation.
Methods
Verify Leave-one-out assumption empirically
@@ Methods
Let ${\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}]^{\top}$ be the input data matrix, and ${\mathbf{Y}}=[{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}$ be the matrix of embedding points. We then add one point $\mathbf{x}$} to $\mathbf{X}$} to have the new input data ${\mathbf{X}}_{+}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n},{\mathbf{x}}]^{\top}$. We then run the t-SNE algorithm to obtain the embedding of $\mathbf{X}$$_{+}$ as $\widetilde{\mathbf{y}}_{1},\ldots,\widetilde{\mathbf{y}}_{n},\widetilde{\mathbf{y}}$$^{\top}$. Denoted $\widetilde{\mathbf{y}}_{1},\ldots,\widetilde{\mathbf{y}}_{n}]$ as $\widetilde{{\mathbf{Y}}}$. To verify LOO empirically, we keep track of the difference between $\mathbf{Y}$} and $\widetilde{{\mathbf{Y}}}$:
$\epsilon_{n}=\frac{1}{\|{\mathbf{Y}}\|_{F}}\|{\mathbf{Y}}-\widetilde{\mathbf{Y}}\|_{F}$
@@
We initialize the t-SNE algorithm in the second run by the embedding points we obtain from the first run: when calculating the embedding of $\mathbf{X}$$_{+}$, we use $\mathbf{Y}$} as the initialization for the first $n$ points. This initialization scheme aims to address two issues: (i) the loss function in a neighbor embedding method is invariant to a global rotation and a global shift of all embedding points, so it is reasonable to choose embedding points with an appropriate initialization. (ii) There are potentially multiple local minima of the loss function due to non-convexity. We verify the LOO assumption at a given local minimum (namely $\mathbf{Y}$}) obtained from the first run.
@@
The experiment is conducted with different sample sizes $n$ and with different types of datasets (simulated cluster data, simulated manifold data, real single-cell data, deep learning feature data). The comprehensive results showing the values of $\epsilon_{n}$ under different settings are presented in Supplementary Table S2. We observe that the approximation errors $\epsilon_{n}$ are small and generally decreasing in $n$, which supports our LOO assumption.
Perturbation score
@@
For implementation convenience, our calculation of the perturbation score and the singularity score is based on modifying an input point instead of adding a new input point. According to the LOO assumption, the difference is negligible.
@@
Given an input data matrix ${\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}]^{\top}$ and its embedding matrix ${\mathbf{Y}}=[{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}$, we view ${\mathbf{y}}_{i}$ as the mapping of ${\mathbf{x}}_{i}$ by the partial LOO-map ${\mathbf{f}}_{i}$:
${\mathbf{f}}_{i}({\mathbf{x}})={\rm argmin}_{{\mathbf{y}}\in\mathbb{R}^{2}}L_{i}({\mathbf{y}};{\mathbf{x}}),\quad\text{where}$ (5)
$L_{i}({\mathbf{y}};{\mathbf{x}})=\sum_{k\neq i}\mathcal{L}\Big(w({\mathbf{y}}_{k},{\mathbf{y}});v_{i,k}(\bar{\mathbf{X}})\Big)+Z(\bar{\mathbf{Y}}),$
@@
where $\bar{\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{i-1},{\mathbf{x}},{\mathbf{x}}_{i+1},\ldots,{\mathbf{x}}_{n}]^{\top}$ differs from $\mathbf{X}$} only at the $i$-th input point, and $\bar{\mathbf{Y}}=[{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{i-1},{\mathbf{y}},{\mathbf{y}}_{i+1},\ldots,{\mathbf{y}}_{n}]^{\top}$ has frozen embedding points except for the $i$-th point which is the decision variable in the optimization problem. This partial LOO-map ${\mathbf{f}}_{i}$ is based on perturbing (or modifying) a single input point rather than adding a new point, thus maintaining $n$ points in total. According to the LOO assumption, ${\mathbf{f}}_{i}\approx{\mathbf{f}}$, so we calculate the perturbation score for the $i$-th point based on ${\mathbf{f}}_{i}$.
@@
To assess the susceptibility of ${\mathbf{y}}_{i}$ under moderate perturbations in ${\mathbf{x}}_{i}$, we apply a perturbation of length $\lambda$ in the direction of $\mathbf{e}$} to ${\mathbf{x}}_{i}$ and measure the resulting change in the embedding map determined by the partial LOO-map ${\mathbf{f}}_{i}$. In our implementation, we search the perturbation directions among the first 3 principal directions of the data $\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},{\mathbf{e}}_{3}\}$ and their opposites $\{-{\mathbf{e}}_{1},-{\mathbf{e}}_{2},-{\mathbf{e}}_{3}\}$, and the perturbation length $\lambda$ is specified by the user. In this way, we can define the perturbation score of the $i$-th data point as
$\max_{{\mathbf{e}}\in\{\pm{\mathbf{e}}_{1},\pm{\mathbf{e}}_{2},\pm{\mathbf{e}}_{3}\}}\|{\mathbf{f}}_{i}({\mathbf{x}}_{i}+\lambda{\mathbf{e}})-{\mathbf{y}}_{i}\|_{2}.$ (6)